OpenStudy (anonymous):

What is the solution to the compound inequality -6<= 2x + 6<=6

OpenStudy (anonymous):

Those are less than or equal to right

OpenStudy (anonymous):

-6<=2x+6<=6 -2x<=6+6 -2x<=12 x>=-2 2x+6<=6 2x<=6-6 2x<=0 x<=0 0>=x>=-2